prove that cos 2 (76)+cos 2 (16)-cos(76)cos(16)=3/4.Give exact proof along with formulas

prove that cos2(76)+cos2(16)-cos(76)cos(16)=3/4.Give exact proof along with formulas
 

Grade:11

1 Answers

Bell Cranel
6 years ago
cos^2 (76) +cos^2 (16) -cos76cos16=c2 (76) +c2 (16) -2c76c16 +c76c16=(c76-c16)^2 +c76c16=(-2sin46sin30)^2 +(c^2 (30) -s^2 (46))=s^2 (46) +c^2 (30) -s^2 (46)=c^2 (30)=3/4